On existence of solution for a class of strongly nonlinear integro-differential equation

Wei Wei

Mathematics department of Guizhou University Guiyang, Guizhou, P. R. China 550025

Abstract

In this paper, a general Gronwall’s lemma with transport delay is proved. A new existence and uniqueness result on solutions for nonlinear integro-differential is presented. The equation contains nonlinear monotone operators and nonlinear nonmonotone perturbations.